OpenStudy (anonymous):

Choose the correct equation in point-slope form for the line through the given point with the given slope. (10, –9); m = –2

OpenStudy (wikiemol):

y-y1 = m(x-x1) so your slope is -2 thats m, x1 is 10 and y1 is -9. to put it in slope intercept form you will have to get y alone on one side and distribute the slope so that you end up with something in the form y = mx + b

OpenStudy (anonymous):

so its y-9= -2(x-10)

OpenStudy (wikiemol):

yes

OpenStudy (anonymous):

thank you sooo much really appreciate the help